OpenStudy (anonymous):

Can someone point me in the right direction, please? f(x) is one-to-one. What is an expression for the inverse of g(x) = f(x+c)?

OpenStudy (anonymous):

@fission-mailed hi i can help you

OpenStudy (anonymous):

@Princer_Jones Thank you :) I know that since f(x) is one-to-one, then \[f^{-1}(x)\] exists. But I'm not quite sure where to go from there.

OpenStudy (anonymous):

A mapping f is said to be one-to-one if f(a)=f(b) implies a=b.

OpenStudy (anonymous):

what is g(x) here?

OpenStudy (anonymous):

@fission-mailed

OpenStudy (anonymous):

Hold on trying to figure it out...

OpenStudy (anonymous):

So 1-1 functions have only one y-value for every x-value. And to find an inverse of a function algebraically you swap the x's and y's and solve for y to get the inverse. So I was thinking y = f(x+c) becomes x = f(y+c), in trying to get the inverse of g(x). But this doesn't seem to make sense, I think I've done something wrong here.

OpenStudy (anonymous):

what are you asked to find?? g(x) or g^-1 (x)?

OpenStudy (anonymous):

I am supposed to find an expression for the inverse of g(x) = f(x+c)

OpenStudy (anonymous):

here g(x) is f(x+c)?

OpenStudy (anonymous):

g(x)=f(x+c)???

OpenStudy (anonymous):

Yes

OpenStudy (aum):

\[ g(x) = f(x+c) \\ \text{Let }~h(x) = x + c. ~~ \text{ Then, } \\ g(x) = f(h(x)) \\ g^{-1}(x) = \left( f(h(x)) \right)^{-1} = h^{-1}(f^{-1}(x)) \\ h(x) = x + c \\ h^{-1}(x) = x - c \\ h^{-1}(f^{-1}(x)) = f^{-1}(x) - c \\ g^{-1}(x) = f^{-1}(x) ~-~ c \]

OpenStudy (aum):

The important thing to remember here is the formula for the inverse of composite functions. If p(x) and q(x) are two functions, then: \[ (p \circ q)^{-1}(x) = (q^{-1} \circ p^{-1})(x) \\ \text{ } \\ \text{Another form of the same formula is:}\\ \text{ } \\ \left\{ p(q(x)\right\}^{-1} = q^{-1}(p^{-1}(x)) \]

OpenStudy (anonymous):

Thanks for the help. I didn't know the formula for the inverse of composite functions before, but it makes sense now.